OpenStudy (anonymous):

Find the product of (x + 3)2

OpenStudy (zzr0ck3r):

\[(a+b)^2 = a^2+2ab+b^2\]

OpenStudy (anonymous):

@zzr0ck3r would it be X^2 + 9

OpenStudy (adilalvi):

x^2+9+6x

OpenStudy (zzr0ck3r):

(x+3)^2 = (x+3)(x+3) = x*x + 3x + 3x + 9 = x^2 + 6x + 9

OpenStudy (anonymous):

thanks !!!! @zzr0ck3r